# LLMs Meet Long Video: Advancing Long Video Comprehension with An Interactive Visual Adapter in LLMs

### v1

@@ Abstract
Long video understanding is a significant and ongoing challenge in the intersection of multimedia and artificial intelligence. Employing large language models (LLMs) for comprehending video becomes an emerging and promising method. However, this approach incurs high computational costs due to the extensive array of video tokens, experiences reduced visual clarity as a consequence of token aggregation, and confronts challenges arising from irrelevant visual tokens while answering video-related questions. To alleviate these issues, we present an Interactive Visual Adapter (IVA) within LLMs, designed to enhance interaction with fine-grained visual elements. Specifically, we first transform long videos into temporal video tokens via leveraging a visual encoder alongside a pretrained causal transformer, then feed them into LLMs with the video instructions. Subsequently, we integrated IVA, which contains a lightweight temporal frame selector and a spatial feature interactor, within the internal blocks of LLMs to capture instruction-aware and fine-grained visual signals. Consequently, the proposed video-LLM facilitates a comprehensive understanding of long video content through appropriate long video modeling and precise visual interactions. We conducted extensive experiments on nine video understanding benchmarks and experimental results show that our interactive visual adapter significantly improves the performance of video LLMs on long video QA tasks. Ablation studies further verify the effectiveness of IVA in long and short video understandings.
@@ 1 Introduction
The exponential advancement of the Internet and multimedia technologies has resulted in a significant surge in video content production by individuals and enterprises across various domains. The ability to interpret and extract meaningful content from videos is increasingly vital for meeting human demands and promoting the speed of information dissemination Tang et al. (2023). Therefore, Video Question Answering Yu et al. (2019); Li et al. (2023b); Castro et al. (2022) (Video QA), which allows users to ask about the content of videos through natural language and receive answers derived from their visual and auditory content, attracts tremendous research interest. Recently, large language models (LLMs) OpenAI (2023); Chiang et al. (2023) have demonstrated exceptional efficacy in the domains of human-machine interaction and the handling of extensive contextual information. Capitalizing on these advancements, there is a burgeoning inclination towards integrating LLMs into the realm of video information processing. This approach primarily aims to enhance the interface between users and video content through intelligent question-and-answer sessions.
@@
These LLMs-powered video understanding models Tang et al. (2023); Song et al. (2023) mainly focus on short video modeling and have achieved a successful performance on short video captioning Zhang et al. (2023c), question-answering Jin et al. (2023), and summarization Tang et al. (2023). However, the core challenges of video processing  Xu et al. (2023) stem from the need to efficiently model long video sequences and precisely respond to questions relevant to the video. Generally, using LLMs to handle long-form video often encounters the following hurdles: 1) high computational costs from a multitude of video tokens; 2) reduced visual clarity as a consequence of token aggregation such as employing average or maximum representation pooling for visual frames; 3) irrelevant visual tokens leading to incorrect answers, notably when question-relevant information is embedded within long temporal cues. Hence, previous models struggle to handle long-form videos owing to the constrained input capacity for video tokens and the challenge of distilling question-relevant, fine-grained visual features during generation.
To alleviate these issues, we present a long video comprehension method for LLMs, named Interactive Visual Adapter (IVA) to achieve in-depth interactions between LLMs and video content. Specifically, we first use the pretrained visual encoder to obtain global and fine-grained frame representations. We construct the temporal video tokens by integrating the global features of frames with temporal video embeddings, which are obtained through a pretrained causal transformer. The whole set of temporal video tokens is fed into the LLM to attain a whole understanding of the video content. Additionally, we designed a parameters-sharing Interactive Visual Adapter (IVA) that contains an instruction-aware temporal frames selector and a spatial feature interactor. The selector is used to obtain question-relevant frames based on contextual query embeddings and global encodings of videos. The selected frames are then fed into the spatial interactor to engage with the contextual query embeddings, in which fine-grained representations of frames are used. By doing so, LLMs could achieve in-depth interaction with video content by applying IVA between different layers.
To verify the effectiveness of our method, we conducted extensive experiments on four long video QA and five short video understanding benchmarks. Experimental results indicate that IVA is capable of achieving effective interactions between LLMs and long or short videos. Our contributions are summarized as follows:
  • We analyze the challenges of modeling long videos for LLMs and propose an interactive visual adapter for LLMs to handle long videos. It realizes the in-depth interaction between LLMs and long videos based on efficient video tokens and the IVA mechanism.
@@
  • Experimental results show that LLMs with IVA could achieve powerful performances in understanding long videos. Ablation studies underscore the critical role and effectiveness of IVA, confirming its significant contribution to enhanced performance.
@@ 2 Related Work
Figure 1: The overview of our framework employing LLMs to handle long video. While producing video tokens, we combine the global features and aggregated fine-grained features to represent a frame, allocating two tokens for each frame. The casual transformer is used to capture temporal relationships across frames and its output will be spliced with spatial feature sequence. The IVA will be inserted between blocks of LLMs to incorporate fine-grained visuals based on an understanding of the long video tokens, text instructions, and query tokens.
LLMs for Video Understanding The recent advancement in large language models (LLMs), pre-trained on expansive datasets, has ushered in groundbreaking capabilities in in-context learning Zhang et al. (2023a) and long-form context modeling Lyu et al. (2023). This innovation has paved the way for integrating LLMs with computer vision technologies, exemplified by initiatives like Visual-ChatGPT Wu et al. (2023). These models transcend traditional boundaries by calling vision model APIs Qin et al. (2023), thereby addressing complex problems within the computer vision domain. Integrating language models with video understanding technologies Maaz et al. (2023); Zhang et al. (2023d); Li et al. (2023e); Xu et al. (2023); Song et al. (2023) enhances multimodal understanding, facilitating sophisticated processing and interpretation of the intricate interplay between visual and textual data. They leverage their extensive multimodal knowledge base and nuanced contextual understanding, mirroring a more human-like comprehension of visual content. Moreover, the exploration of LLMs in video understanding tasks Tang et al. (2023) represents a significant stride towards harnessing their potential in analyzing and reasoning about visual data.
Multimodal Instruction Tuning for LLMs Recent advancements have significantly enhanced the performance of instruction-tuned, text-only large language models (LLMs) Ouyang et al. (2022); Muennighoff et al. (2022); Li et al. (2023g) on various NLP tasks and human-machine interaction scenarios. SFT LLMs have demonstrated remarkable capabilities in these areas. Building on this foundation, researchers are now exploring the integration of multimodal instruction data to further refine pre-trained LLMs, aiming to elevate their multimodal human-machine interaction competencies. For instance, a study Liu et al. (2023) employed GPT-4 to generate multimodal instruction data, which was then used to fine-tune the language model LLaMA on a synthetic dataset designed for instruction following. Similarly, another research Zhu et al. (2023) effort constructed a well-aligned multimodal instruction-following dataset to fine-tune Vicuna, an instruction-tuned language model, which showed superior performance in open-domain multimodal dialogues. Furthermore, a lightweight adaptation method Zhang et al. (2023d); Li et al. (2023h) was introduced to efficiently convert LLaMA into an instruction-following model, showcasing the potential for streamlined model enhancement. In this paper, we also explore the lightweight adapter to help LLMs understand long videos.
@@ 3.1 Overview
Our work primarily introduces an interactive visual adapter for LLMs to handle long videos and answer relevant questions. The overview of workflow is shown in Figure 1. Specifically, given a video $V$, we first extract frames to obtain the whole sequence frame representations $\mathbf{h}_{V}=(\mathbf{h}_{I_{1}},...,\mathbf{h}_{I_{k}},...,\mathbf{h}_{I_{N}})$ via the pretrained image encoder, where $\mathbf{h}_{I_{k}}=(h_{g}^{I_{k}},h_{1}^{I_{k}}...,h_{576}^{I_{k}})$ refers to the representations of $k$ th frame and $N$ is the total number of extracted frames. Then, we use a casual transformer to acquire temporal video embeddings from the aggregated spatial representation. The overall video tokens are formed by merging temporal video embeddings and global spatial features $[h_{g}^{I_{1}},h_{g}^{I_{k}},...,h_{g}^{I_{N}}]$, where each frame is represented by two tokens. To enhance the capability of LLMs in leveraging fine-grained visual details from videos, we have developed an Interactive Visual Adapter (IVA) that is integrated into the blocks of LLM. This integration allows LLMs to comprehend the entirety of long videos through efficient video tokens while simultaneously capturing fine-grained visual information facilitated by the IVA.
@@ 3.2 Producing Video Tokens
We elaborate on the detailed process employed to produce efficient tokens for long videos, characterized by the extraction of one frame per second. First, we use the self-weighted calculation on the fine-grained feature $\mathbf{h}_{f}^{k}=(h_{1},...,h_{576})$ of a frame ($k$ th) to obtain its overall representation, which will be fed into the following casual transformer. This calculation process for the $k$ th frame is given in the following Eq. 1:
$\begin{array}[]{cc}s_{f}^{k}=Softmax(\mathbf{W}(\mathbf{h}_{f}^{k})+\mathbf{b}),\\ \mathbf{h}_{t}^{k}=s_{f}\mathbf{h}_{f}^{k},\end{array}$ (1)
@@
Casual Transformer is employed to acquire the temporal video embeddings. Specifically, we use a four-layer transformer to facilitate interaction across frames, where a frame only attends to its previous ones. Take the first layer as an example, the specific operation of the casual transformer is presented in Eq. 2:
$\begin{array}[]{cc}\mathbf{h}_{s}=SelfAtten(LayerN(\mathbf{h}_{f}),W^{Mask})+\mathbf{h}_{f}\\ \mathbf{h}_{s}=LayerN(\mathbf{h}_{s}),\\ \mathbf{h}_{o}^{1}=MLP(\mathbf{h}_{s})\end{array}$ (2)
where $SelfAtten$ and $LayerN$ are the self-attention calculation and the feature normalization. The top output of the casual transformer will be projected into the language model by a linear layer, which is spliced with the global features $\mathbf{h}_{g}^{V}=(h_{g}^{I_{1}},h_{g}^{I_{k}},...,h_{g}^{I_{N}})$. These global features of frames will be transferred into language models via a learnable MLP. We denote the final spliced feature to $\mathbf{h}_{V}=(h_{V}^{1},h_{V}^{2},...,h_{V}^{2N})$.
@@ 3.3 Interactive Visual Adapter
$\begin{array}[]{cc}\mathbf{h}^{S}_{q}=W^{q}\mathbf{h}^{i}_{d}+b^{q},\\ \mathbf{h}^{S}_{k}=W^{k}\mathbf{h}_{g}^{V}+b^{k},\\ \mathbf{M}^{S}=\mathbf{h}^{S}_{q}(\mathbf{h}^{S}_{k})^{T},\\ \mathbf{h}^{S}=Softmax(\mathbf{M}^{S}/\tau)Trans([\mathbf{h}_{I_{1}}^{f},...,\mathbf{h}_{I_{N}}^{f}]),\end{array}$ (3)
@@
$\begin{array}[]{cc}\mathbf{h}^{I}_{q}=W^{1}\mathbf{h}^{i}_{d}+b^{1},\\ \mathbf{h}^{I}_{k}=W^{2}\mathbf{h}^{S}+b^{2},\\ \mathbf{M}^{I}=\mathbf{h}^{I}_{q}(\mathbf{h}^{I}_{k})^{T},\\ \mathbf{h}^{S}_{c}=Softmax(\mathbf{M}^{I})(W^{3}\mathbf{h}^{S}+b^{3}),\\ \mathbf{h}^{S}=MLP(\mathbf{h}^{S}_{c})+\mathbf{h}^{S}_{c}\end{array}$ (4)
@@ 3.4 Training
Stage 1: Pretraining. To endow video tokens with meaningful representation, we first train the casual transformer, linear layers, and other learnable parameters during video tokens production, on massive video-caption pairs from WebVid, a total of 703k video-caption pairs. We freeze the other parameters of the overall model during this process and do not introduce the IVA module.
@@
where $r$ represents the round count. As shown in Eq. 5, when $r>$ 1, we concatenate the conversations from all previous rounds with the current instruction as the input for this round. The training objective remains the same as the previous stage. After this stage, the model will be capable of generating corresponding responses based on various instructions.
@@
Video-LLaMA Zhang et al. (2023b) 12.4 1.1 55.8 2.9 35.8 2.3 63.2 3.2
@@
Video-ChatGPT Maaz et al. (2023) 35.2 2.7 57.5 3.3 33.9 2.6 58.0 3.3
Baseline (w/o IVA) 40.8 3.0 46.8 3.0 31.7 2.3 52.2 3.1
@@
Table 1: Comparison between different methods on 4 zero-shot long video QA datasets. LLM with IVA achieves best performance on long videos compared to baselines and strong video LLMs. “LQ” refers to the length of query tokens and “NI” represents the number of interactions between LLMs and IVA. “-272K” indicates that we introduce additional training data of long video datasets like LifeQA and Social-IQ based on the original short video data.
@@ 4.1 Data sets
While training the casual transformer, we utilize 702 thousand video-text pairs derived from Valley Luo et al. (2023), sourced from WebVid Bain et al. (2021). For the Video Instruction Tuning stage, we first collect instructional datasets from three sources. This includes a 100K video-text instruction dataset from Video-ChatGPT Muhammad Maaz and Khan (2023), a 36K short video-text instruction dataset from Valley-Instruct-73k Luo et al. (2023), and a 34K multiple-choice QA dataset from NExT-QA Xiao et al. (2021).
Additionally, we assessed the generalization of IVA using long and short video benchmarks. Long video benchmarks typically are characterized by videos exceeding one minute in duration. We evaluated our model using four prominent long video evaluation benchmarks: ActivityNet-QA Yu et al. (2019), Social-IQ 2.0 Wilf et al. (2023), LifeQA Castro et al. (2020), WildQA Castro et al. (2022). For short video benchmarks, the duration of the videos is often measured in seconds. We evaluated our model against three notable short video evaluation benchmarks: MSVD-QA Xu et al. (2017), MSRVTT-QA Xu et al. (2017), and SEED-Bench Li et al. (2023c).

### v2

# LLMs Meet Long Video: Advancing Long Video Question Answering with An Interactive Visual Adapter in LLMs

@@ Abstract
Long video understanding is a significant and ongoing challenge in the intersection of multimedia and artificial intelligence. Employing large language models (LLMs) for comprehending video becomes an emerging and promising method. However, this approach incurs high computational costs due to the extensive array of video tokens, experiences reduced visual clarity as a consequence of token aggregation, and confronts challenges arising from irrelevant visual tokens while answering video-related questions. To alleviate these issues, we present an Interactive Visual Adapter (IVA) within LLMs, designed to enhance interaction with fine-grained visual elements. Specifically, we first transform long videos into temporal video tokens via leveraging a visual encoder alongside a pretrained causal transformer, then feed them into LLMs with the video instructions. Subsequently, we integrated IVA, which contains a lightweight temporal frame selector and a spatial feature interactor, within the internal blocks of LLMs to capture instruction-aware and fine-grained visual signals. Consequently, the proposed video-LLM facilitates a comprehensive understanding of long video content through appropriate long video modelling and precise visual interactions. We conduct extensive experiments on nine video understanding benchmarks and experimental results show that our interactive visual adapter significantly improves the performance of video LLMs on long video QA tasks. Ablation studies further verify the effectiveness of IVA in long and short video understandings.
@@ 1 Introduction
The exponential advancement of the Internet and multimedia technologies has resulted in a significant surge in video content production by individuals and enterprises across various domains. The ability to interpret and extract meaningful content from videos is increasingly vital for meeting human demands and promoting the speed of information dissemination Tang et al. (2023). Therefore, Video Question Answering Yu et al. (2019); Li et al. (2023b); Castro et al. (2022) (Video QA), which allows users to ask about the content of videos through natural language and receive answers derived from their visual and auditory content, attracts tremendous research interest. Recently, large language models (LLMs) OpenAI (2023); Chiang et al. (2023) have demonstrated exceptional efficacy in the domains of human-machine interaction and the handling of extensive contextual information. Capitalizing on these advancements, there is a burgeoning inclination towards integrating LLMs into the realm of video information processing. This approach primarily aims to enhance the interface between users and video content through intelligent question-and-answer (QA) sessions.
@@
These LLMs-powered video understanding models Tang et al. (2023); Song et al. (2023) mainly focus on short video modelling and have achieved a successful performance on short video captioning Zhang et al. (2023c), question-answering Jin et al. (2023), and summarization Tang et al. (2023). However, the core challenges of video processing  Xu et al. (2023) stem from the need to efficiently model long video sequences and precisely respond to questions relevant to the video. Generally, using LLMs to handle long-form video often encounters the following hurdles: 1) high computational costs from a multitude of video tokens; 2) reduced visual clarity as a consequence of token aggregation such as employing average or maximum representation pooling for visual frames; 3) irrelevant visual tokens leading to incorrect answers, notably when question-relevant information is embedded within long temporal cues. Hence, previous models struggle to handle long-form videos owing to the constrained input capacity for video tokens and the challenge of distilling question-relevant, fine-grained visual features during generation.
To alleviate these issues, we present a long video comprehension method for LLMs, named Interactive Visual Adapter (IVA) to achieve in-depth interactions between LLMs and video content. Specifically, we first use the pretrained visual encoder to obtain global and fine-grained frame representations. We construct the temporal video tokens by integrating the global features of frames with temporal video embeddings, which are obtained through a pretrained causal transformer. The whole set of temporal video tokens is fed into the LLM to attain a whole understanding of the video content. Additionally, we design a parameters-sharing Interactive Visual Adapter (IVA) that contains an instruction-aware temporal frames selector and a spatial feature interactor. The selector is used to obtain question-relevant frames based on contextual query embeddings and global encodings of videos. The selected frames are then fed into the spatial interactor to engage with the contextual query embeddings, in which fine-grained representations of frames are used. By doing so, LLMs could achieve in-depth interaction with video content by applying IVA between different layers.
To verify the effectiveness of our method, we conduct extensive experiments on four long video QA and five short video understanding benchmarks. Experimental results indicate that IVA is capable of achieving effective interactions between LLMs and long or short videos. Our contributions are summarized as follows:
  • We analyze the challenges of modelling long videos for LLMs and propose an interactive visual adapter for LLMs to handle long videos. It realizes the in-depth interaction between LLMs and long videos based on efficient video tokens and the IVA mechanism.
@@
  • Experimental results show that LLMs with IVA could achieve powerful performances in long video QA. Ablation studies underscore the critical role and effectiveness of IVA, confirming its significant contribution to enhanced performance.
Figure 1: The overview of our framework employing LLMs to handle long video. While producing video tokens, we combine the global features and aggregated fine-grained features to represent a frame, allocating two tokens for each frame. The causal transformer is used to capture temporal relationships across frames and its output will be spliced with spatial feature sequence. The IVA will be inserted between blocks of LLMs to incorporate fine-grained visuals based on an understanding of the long video tokens, text instructions, and query tokens.
@@ 2 Related Work
LLMs for Video Understanding The recent advancement in large language models (LLMs), pre-trained on expansive datasets, has ushered in groundbreaking capabilities in in-context learning Zhang et al. (2023a) and long-form context modeling Lyu et al. (2023). This innovation has paved the way for integrating LLMs with computer vision technologies, exemplified by initiatives like Visual-ChatGPT Wu et al. (2023). These models transcend traditional boundaries by calling vision model APIs Qin et al. (2023), thereby addressing complex problems within the computer vision domain. Integrating language models with video understanding technologies Maaz et al. (2023); Zhang et al. (2023d); Li et al. (2023e); Xu et al. (2023); Song et al. (2023) enhances multimodal understanding, facilitating sophisticated processing and interpretation of the intricate interplay between visual and textual data Ouyang et al. (2022); Liu et al. (2023); Muennighoff et al. (2022); Li et al. (2023g); Zhu et al. (2023); Zhang et al. (2023d). They leverage their extensive multimodal knowledge base and nuanced contextual understanding, mirroring a more human-like comprehension of video. Moreover, the exploration of LLMs in video understanding tasks Tang et al. (2023) represents a significant stride in analyzing and reasoning about visual data.
@@ 3.1 Overview
Our work primarily introduces an interactive visual adapter for LLMs to handle long videos and answer relevant questions. The overview of workflow is shown in Figure 1. Specifically, given a video $V$, we first extract frames to obtain the whole sequence frame representations $\mathbf{h}_{V}=(\mathbf{h}_{I_{1}},...,\mathbf{h}_{I_{k}},...,\mathbf{h}_{I_{N}})$ via the pretrained image encoder, where $\mathbf{h}_{I_{k}}=(h_{g}^{I_{k}},h_{1}^{I_{k}}...,h_{576}^{I_{k}})$ refers to the representations of $k$ th frame and $N$ is the total number of extracted frames. Then, we use a causal transformer to acquire temporal video embeddings from the aggregated spatial representation. The overall video tokens are formed by merging temporal video embeddings and global spatial features $[h_{g}^{I_{1}},h_{g}^{I_{k}},...,h_{g}^{I_{N}}]$, where each frame is represented by two tokens. To enhance the capability of LLMs in leveraging fine-grained visual details from videos, we have developed an Interactive Visual Adapter (IVA) that is integrated into the blocks of LLM. This integration allows LLMs to comprehend the entirety of long videos through efficient video tokens while simultaneously capturing fine-grained visual information facilitated by the IVA.
@@ 3.2 Producing Video Tokens
We elaborate on the detailed process employed to produce efficient tokens for long videos, characterized by the extraction of one frame per second. First, we use the self-weighted calculation on the fine-grained feature $\mathbf{h}_{f}^{k}=(h_{1},...,h_{576})$ of a frame ($k$ th) to obtain its overall representation, which will be fed into the following causal transformer. This calculation process for the $k$ th frame is given in the following Eq. 1:
$\begin{array}[]{cc}s_{f}^{k}=Softmax(\mathbf{W}(\mathbf{h}_{f}^{k})+\mathbf{b}),\vspace{1.0ex}\\ \mathbf{h}_{t}^{k}=s_{f}\mathbf{h}_{f}^{k},\end{array}$ (1)
@@
Causal Transformer is employed to acquire the temporal video embeddings. Specifically, we use a four-layer transformer to facilitate interaction across frames, where a frame only attends to its previous ones. Take the first layer as an example, the specific operation of the causal transformer is presented in Eq. 2:
$\begin{array}[]{cc}\mathbf{h}_{s}=SelfAtten(LayerN(\mathbf{h}_{f}),W^{Mask})+\mathbf{h}_{f}\vspace{1.0ex}\\ \mathbf{h}_{s}=LayerN(\mathbf{h}_{s}),\vspace{1.0ex}\\ \mathbf{h}_{o}^{1}=MLP(\mathbf{h}_{s})\end{array}$ (2)
where $SelfAtten$ and $LayerN$ are the self-attention calculation and the feature normalization. The top output of the causal transformer will be projected into the language model by a linear layer, which is spliced with the global features $\mathbf{h}_{g}^{V}=(h_{g}^{I_{1}},h_{g}^{I_{k}},...,h_{g}^{I_{N}})$. These global features of frames will be transferred into language models via a learnable MLP. We denote the final spliced feature to $\mathbf{h}_{V}=(h_{V}^{1},h_{V}^{2},...,h_{V}^{2N})$.
@@ 3.3 Interactive Visual Adapter
$\begin{array}[]{cc}\mathbf{h}^{S}_{q}=W^{q}\mathbf{h}^{i}_{d}+b^{q},\vspace{1.0ex}\\ \mathbf{h}^{S}_{k}=W^{k}\mathbf{h}_{g}^{V}+b^{k},\vspace{1.0ex}\\ \mathbf{M}^{S}=\mathbf{h}^{S}_{q}(\mathbf{h}^{S}_{k})^{T},\vspace{1.0ex}\\ \mathbf{h}^{S}=Softmax(\mathbf{M}^{S}/\tau)Trans([\mathbf{h}_{I_{1}}^{f},...,\mathbf{h}_{I_{N}}^{f}]),\end{array}$ (3)
@@
$\begin{array}[]{cc}\mathbf{h}^{I}_{q}=W^{1}\mathbf{h}^{i}_{d}+b^{1},\vspace{1.0ex}\\ \mathbf{h}^{I}_{k}=W^{2}\mathbf{h}^{S}+b^{2},\vspace{1.0ex}\\ \mathbf{M}^{I}=\mathbf{h}^{I}_{q}(\mathbf{h}^{I}_{k})^{T},\vspace{1.0ex}\\ \mathbf{h}^{S}_{c}=Softmax(\mathbf{M}^{I})(W^{3}\mathbf{h}^{S}+b^{3}),\vspace{1.0ex}\\ \mathbf{h}^{S}=MLP(\mathbf{h}^{S}_{c})+\mathbf{h}^{S}_{c}\end{array}$ (4)
@@ 3.4 Training
Stage 1: Pretraining. To endow video tokens with meaningful representation, we first train the causal transformer, linear layers, and other learnable parameters during video tokens production, on massive video-caption pairs from WebVid, a total of 703k video-caption pairs. We freeze the other parameters of the overall model during this process and do not introduce the IVA module.
@@
where $r$ represents the round count. As shown in Eq. 5, when $r>$ 1, we concatenate the conversations from all previous rounds with the current instruction as the input for this round. The training objective remains the same as the previous stage. After this stage, the model can generate corresponding responses for various instructions.
@@
Video-LLaMA Zhang et al. (2023b) 12.4 1.1 48.2 2.8 28.8 2.3 57.5 3.2
@@
Video-ChatGPT Maaz et al. (2023) 35.2 2.7 51.6 3.2 31.2 2.5 54.9 3.3
Baseline (w/o IVA) w/o Spatial Token 38.4 3.0 49.8 3.1 26.3 2.2 51.4 3.0
Baseline (w/o IVA) w/o Causal Token 38.2 2.9 49.9 3.1 30.6 2.4 51.7 3.1
Baseline (w/o IVA) 40.8 3.0 53.0 3.3 30.9 2.5 54.4 3.2
@@
Table 1: Comparison between different methods on 4 long video QA datasets. LLM with IVA achieves the best performance on long videos compared to baselines and strong video LLMs. “LQ” refers to the length of query tokens and “NI” represents the number of interactions between LLMs and IVA. “-272K” indicates that we introduce additional training data of long video datasets like LifeQA and Social-IQ based on the original short video data.
@@ 4.1 Data sets
While training the causal transformer, we utilize 702 thousand video-text pairs derived from Valley Luo et al. (2023), sourced from WebVid Bain et al. (2021). During the instruction tuning stage, we collect instructional datasets from three sources: a 100K video-text instruction dataset from Video-ChatGPT Muhammad Maaz and Khan (2023), a 36K short video-text instruction dataset from Valley-Instruct-73k Luo et al. (2023), and a 34K multiple-choice QA dataset from NExT-QA Xiao et al. (2021). Additionally, we assess the generalization of IVA using long and short video benchmarks. Long video benchmarks typically are characterized by videos exceeding one minute in duration. We evaluated our model using four prominent long video evaluation benchmarks: ActivityNet-QA Yu et al. (2019), Social-IQ 2.0 Wilf et al. (2023), LifeQA Castro et al. (2020), WildQA Castro et al. (2022). For short video benchmarks, the duration of the videos is often measured in several seconds. We evaluate our model against three notable short video evaluation benchmarks: MSVD-QA Xu et al. (2017), MSRVTT-QA Xu et al. (2017), and SEED-Bench Li et al. (2023c).